Multiply and reduce to lowest form (if possible):
(3/8) × (6/4)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then,
= (3 × 6)/ (8 × 4)
= (3 × 3)/ (4 × 4)
= (9/16)

Thanks